OpenStudy (anonymous):

Given the function f(x) = 3(x − 3) 2 + 2, indicate the shifts that will effect the location of the vertex and explain what effect they will have. Use complete sentences. f(x + 4) f(x) + 4 f(4x) 4•f(x)

OpenStudy (anonymous):

dam this is confusing

OpenStudy (anonymous):

well thank you for the statement but I really need help.

OpenStudy (anonymous):

i've been trying to solve it for 15 minutes

OpenStudy (anonymous):

f(x) = 3(x − 3)^2 + 2 is like f(x) = a(x-h)^2 + k where h, k is the vertex of the parabola, so the vertext of the parabola is (3, 2) so, if we have f(x+4) then f(x+4) = 3((x+4) − 3)^2 + 2 f(x+4) = 3(x+4 − 3)^2 + 2 f(x+4) = 3(x+1)^2 + 2 f(x+4) = 3(x-(-1))^2 + 2 so, this now has a vertex of (-1, 2), so the f(x+4) shifted the vertex to the left 4 units. making sense?

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (anonymous):

^_^